OpenStudy (anonymous):

Henry went fishing and caught 4 fish weighing 3 1/2, 5 1/4, 2 1/2, and 6 1/4. What is the average weight of the 4 fish henry caught?

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (angela210793):

u add all those and then divide them by 4

OpenStudy (anonymous):

17.5/4

OpenStudy (anonymous):

well they only give the choices of 5 1/4, 5, 4 3/8, or 4 1/8

OpenStudy (anonymous):

i think brackett forgot to divide by 4

OpenStudy (anonymous):

mixed numbers are a drag

OpenStudy (anonymous):

lol i did it and got 4 2/8 but its not an answer choice

OpenStudy (anonymous):

oh he wrote \[\frac{17.5}{4}\]

OpenStudy (angela210793):

7/2+21/4+5/2+25/4=12/2+46/4=6+11.5 =17.5 then the average=17.5/4

OpenStudy (anonymous):

think it is 3/8 not 2/8

OpenStudy (anonymous):

\[\frac{17.5}{4}=\frac{35}{8}=4\tfrac{3}{8}\]

OpenStudy (anonymous):

ok ty

OpenStudy (anonymous):

whew!